OpenStudy (anonymous):

I NEED HELP WILL FAN AND MEDAL! Solve for n. Round to the tenths place, if necessary 14/21.2=n/5

OpenStudy (anonymous):

This? \[\frac{ 14 }{ 21.2 }=\frac{ n }{ 5 }\]

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

no, I was asking if that's your equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. so it's dividing n by 5. To you undo that, multiply both sides by 5.

OpenStudy (anonymous):

\[\frac{ 5*14 }{ 21.2 }=n\]

OpenStudy (anonymous):

k. Thanks I gtg.